Calculate the SVD of the matrix 3 2 A 2 3 2 -2 Based on the SVD results, express range(A) and null(A)
Answer #1

2 So 丈 mit veeter Haling an 3 2/3 2 Tha ea - 2 2 2. -2 -1/4 eign er isan nelols we kol v,= Fe A=2s vsirg Similas 2/3 1/3 To c3 2 A = 2 2 tat A -U v unlayy a Triangular nat Surh TO O 2 A 3 2. 2 -2 The chasatois petyia A*A dt AA-) = 4-aa-1) 2. 3 4 > +all A) U2 vt- o at A o lesh O => C .lA)=fol To indorge (A) Colums U Z |3AR )- Spo Hence ang (A 12 suods